OpenStudy (luigi0210):

Integrate:

OpenStudy (luigi0210):

\[\LARGE \int x^2~\sqrt{2+x}~dx\]

OpenStudy (kainui):

2+x=u it's kinda tricky.

OpenStudy (anonymous):

Substitute \(u=x+2\), then \(du=dx\) and \(x^2=(u-2)^2\).

OpenStudy (kainui):

I wonder if you could solve this by making the substitution \[x=2\tan^2 \theta\] lol

OpenStudy (luigi0210):

This is the type of u-sub that I never understood. So we end up with \[\LARGE \int (u-2)^2 ~\sqrt{u}~du\]

OpenStudy (kainui):

Yes, does it make sense how we get there though? \[\int\limits x^2 \sqrt{2+x} \ dx\] \[u=2+x\]\[u-2=x\]\[\frac{du}{dx}=1\]\[du=dx\] \[\int\limits(u-2)^2\sqrt{u} \ du\]\[\int\limits (u^2-2u+4)u^{1/2}\ du\]\[\int\limits u^{5/2}-2u^{3/2}+4u^{1/2}\ du\]

OpenStudy (kainui):

I totally squared that wrong, should be (u^2-4u+u) lol.

OpenStudy (luigi0210):

Ohh, haha, I see what you were trying to do tho, thank you for the explanation, really helped! :)

OpenStudy (kainui):

Glad I could help!

sammixboo (sammixboo):

Still going... WOOHOO